OpenStudy (tcp2315):

so didnt pay attention is first calc class a week ago and left the hmwk until tonight.. who can help me "evaluate" the following...

OpenStudy (anonymous):

lol good luck

OpenStudy (tcp2315):

\[\lim_{x \rightarrow 2}(x ^{2}+4x+1)\]

OpenStudy (mertsj):

13

OpenStudy (tcp2315):

so then \[\lim_{h \rightarrow 0}((h-5)^{2}-25) \div h\] is 0

OpenStudy (anonymous):

unfortunately no, since the result of that is undefined (0/0) you have to do other stuff to compute the limit (remember a limit doesn't care if the function is defined or not)

OpenStudy (tcp2315):

so is that when i have to find the "hole"?

OpenStudy (anonymous):

do you know l'hopitals rule?

OpenStudy (anonymous):

thats probably what went on in your lecture this week!

OpenStudy (anonymous):

if hes in the first part of calculus they probably want him to do that w/ algebra and not lhopitals rule (that's what they made me do)

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{h^2-10h+25-25}{h}=\lim_{h \rightarrow 0}\frac{h^2-10h}{h}=\lim_{h \rightarrow 0}(h-10)\]

OpenStudy (anonymous):

yeah that

OpenStudy (anonymous):

yes this is newtons approach to taking derivatives! Very good lesson! Calculus I good lord that was long ago!

OpenStudy (tcp2315):

haha high school calculus, payed attention next calss. that concept was unbelievably easy i must have looked stupid asking that hahahah. wow.

myininaya (myininaya):

There are no stupid questions.